# Prompt-Based Bias Calibration for Better Zero/Few-Shot Learning of Language Models

### v1

@@ Abstract
Prompt learning is susceptible to intrinsic bias present in pre-trained language models (LMs), resulting in sub-optimal performance of prompt-based zero/few-shot learning. In this work, we propose a null-input prompting method to calibrate intrinsic bias encoded in pre-trained LMs. Different from prior efforts that address intrinsic bias primarily for social fairness and often involve excessive computational cost, our objective is to explore enhancing LMs’ performance in downstream zero/few-shot learning while emphasizing the efficiency of intrinsic bias calibration. Specifically, we leverage a diverse set of auto-selected null-meaning inputs generated from GPT-4 to prompt pre-trained LMs for intrinsic bias probing. Utilizing the bias-reflected probability distribution, we formulate a distribution disparity loss for bias calibration, where we exclusively update bias parameters ( of total parameters) of LMs towards equal probability distribution. Experimental results show that the calibration promotes an equitable starting point for LMs while preserving language modeling abilities. Across a wide range of datasets, including sentiment analysis and topic classification, our method significantly improves zero/few-shot learning performance of LMs for both in-context learning and prompt-based fine-tuning (on average and , respectively).
@@ 1 Introduction
The advent of GPT models Radford et al. (2019); Brown et al. (2020) has catalyzed the transformative prompt-learning paradigm. The innovative approach of "pre-train, prompt, and predict" Schick and Schütze (2021a); Liu et al. (2023) facilitates fast adaptation of pre-trained language models (LMs) in learning various tasks and empowering LMs’ strong zero/few-shot learning abilities Schick and Schütze (2021b); Gao et al. (2021).
@@
Due to the susceptibility to bias ingrained in pre-trained LMs, prompt learning tends to make biased predictions toward some specific answers, thereby impacting the performance of prompt-based zero/few-shot learning Zhao et al. (2021); Han et al. (2023). To mitigate this issue and improve LM performance, Zhao et al. (2021) and Holtzman et al. (2022) propose to reweigh LM output probabilities. Han et al. (2023) explores calibrating decision boundaries. While these research has demonstrated substantial improvements, they are primarily designed for in-context learning with frozen pre-trained LMs, leading to two main limitations: (1) They may be not effective in task-specific fine-tuning scenario Jian et al. (2022). Note, however, prompt-based fine-tuning has shown performance improvements over in-context learning Gao et al. (2021); Logan IV et al. (2022). It is particularly important for relatively small-sized LMs. (2) The intrinsic bias encoded in pre-trained LMs persists since these research focuses on output calibration and does not modify LMs.
@@
To address these limitations, we investigate the potential for enhancing the performance of LMs as zero/few-shot learners in classification tasks by calibrating intrinsic bias of pre-trained LMs. This exploration extends to various prompt-learning scenarios: in-context learning and prompt-based fine-tuning. Prior approaches to mitigate intrinsic bias primarily focus on achieving social fairness, and often require laborious corpora augmentation and costly re-training Huang et al. (2020); Kaneko and Bollegala (2021); Solaiman and Dennison (2021); Li et al. (2023). To improve efficiency in both data generation and model updates, we propose leveraging auto-generated null-meaning inputs to prompt LMs for intrinsic bias probing, and subsequently updating only bias parameters BLM of LMs for bias calibration. Our motivation stems from the expectation that bias-calibrated models should produce uniform probabilities across all categories if the input in a prompt delivers null information Zhao et al. (2021). BLM functions as offsets in neural networks, and strategically updating only BLM could potentially counteract intrinsic bias of pre-trained models, achieving higher efficiency (updating parameters of entire LM). The approach promotes an equitable starting point, and we expect that the light model updates preserve pre-trained models’ language understanding abilities while maintaining the focus on bias calibration, ultimately making LMs better zero/few-shot learners.
@@
The pipeline of our calibration method is illustrated in Figure 1. We use Masked LMs (RoBERTa Liu et al., 2019) for zero/few-shot learning since they generally produce competitive performance in classification tasks and their moderate size facilitates combining prompting with fine-tuning Gao et al. (2021); Liu et al. (2023). First, we utilize GPT-4 API to automatically generate diverse null-meaning inputs including symbols, words, phrases, and sentences. This generation process is downstream task-agnostic. By concatenating each null-meaning input with an answer format ans aligned with the downstream task, we construct null-input prompts (similar to Zhao et al., 2021), e.g., "An empty sentence. It is about <mask>.". For better cohesive integration of the "null" information into the prompts, we additionally devise a filtering strategy to select , to which the answer format ans exhibits relatively strong Next Sentence Prediction (NSP) correlation Devlin et al. (2019). Next, we update BLM with null-input prompts to calibrate intrinsic bias. Given the absence of task-relevant information in these prompts, the anticipated outcome in the parameter updating process is a convergence towards equal output probabilities for each label word. We formulate a customized Kullback–Leibler (KL) divergence loss Kullback and Leibler (1951) for gradient descent on BLM to minimize the distribution disparity. Finally, bias-calibrated LMs are applied in downstream prompt-based zero/few-shot learning following Gao et al. (2021).
@@
Extensive experiments on eight classification datasets with four prompt-learning approaches show that our method significantly improves LMs’ zero/few-shot performance, and outperforms output-calibration methods.
2 Preliminaries
Impact of intrinsic bias on downstream LM performance. Intrinsic bias in pre-trained LMs stems from imbalances present in extensive pre-training corpora. Higher frequency of specific terms in those corpora could lead to common token bias Zhao et al. (2021). Additionally, frequent co-occurrence of certain terms with specific sentiment in pre-training could introduce association bias111For example, in a restaurant review “I had roast chicken and a salad.”, RoBERTa model classifies the sentiment for “roast chicken” as positive, while the true label is neutral. This may arise from the association of ”roast chicken” with positive sentiment words prevalent in pre-training corpora. Cao et al. (2022). Because of those intrinsic bias, prompt-based predictions by pre-trained LMs are prone to bias towards some specific answers, resulting in sub-optimal performance in downstream tasks Zhao et al. (2021); Han et al. (2023).
Mitigating strategies from related work. Research has focused on counteracting the bias solely at the output prediction stage, without modifying pre-trained LMs. For example, Zhao et al. (2021) introduces contextual calibration and Holtzman et al. (2022) presents Domain Conditional Pointwise Mutual Information to reweigh answer scores. Min et al. (2022) explores computing the probability of the input conditioned on the label. Han et al. (2023) proposes to calibrate decision boundaries. However, these studies mainly demonstrate their effectiveness for in-context learning using frozen pre-trained LMs, without addressing the intrinsic bias encoded in the LMs. Other research on mitigating intrinsic bias primarily targets removing social bias Dinan et al. (2020); Huang et al. (2020); Cheng et al. (2021); Zhou et al. (2023), often employing costly data augmentation and re-training, and as a by-product, degrades language modeling abilities Garimella et al. (2021); Meade et al. (2022).
Efficiently calibrating intrinsic bias in pre-trained LMs for enhancing downstream zero/few-shot prompt learning performance is an open research problem. In this work, we introduce a parameter-efficient intrinsic-bias calibration method leveraging automatically constructed null-input prompts. We demonstrate its effectiveness of making LMs better zero/few-shot learners for both in-context learning and prompt-based fine-tuning.
@@ 3.1 Task Formulation
where denotes uniform probability distribution and is simplified as .
@@ 3.2 Update Only Bias Parameters
While intrinsic bias may be encoded across various parts of pre-trained LMs, one question arises: is it essential to update the entire model, or is there a more efficient alternative that can achieve comparable effectiveness in intrinsic bias calibration? We propose to only update bias parameters BLM, with the following rationale: (i) BLM constitutes less than of total LM parameters, offering significant memory and computation cost saving compared to updating entire LM. (ii) Weight parameters WLM222WLM also includes embedding parameters in our context. may carry crucial pre-existing knowledge for language modeling, which risks impairment with a full model update Meade et al. (2022). BLM, often overlooked in LM research, serves as offsets in DNN layers. Strategic updates may counteract intrinsic bias while potentially preserving language modeling abilities. (iii) Empirical research on efficient fine-tuning has demonstrated the important role of bias parameters in LMs Ben Zaken et al. (2022); Logan IV et al. (2022).
@@ 3.3 Early Stopping of Calibration
We aim to obtain LM with improved zero/few-shot performance at the calibration stopping point. An overly calibrated model may simply produce uniform probability predictions regardless of input information, deviating from our intended objective. We develop different early stopping strategies depending on whether the downstream task is zero-shot or few-shot.
For zero-shot downstream tasks. Determining the calibration stopping point for optimal zero-shot learning performance is challenging due to the absence of labeled data for validation during calibration. To discern the patterns of a good stopping point, we first conduct empirical experiments by validating LM zero-shot performance on the entire test dataset after each calibration batch (consisting of null-meaning inputs). As shown in Figure 2, we observe that model performance has steep increases in the first one/few calibration batches333We experimented with smaller calibration learning rates and observed consistent less improvement of LM performance. with low variance, and then starts to degrade and becomes unstable. The low performance and instability at the calibration tail confirm our assumption on the detrimental effects of excessive calibration on LM’s modeling abilities. Notably, calibration with only one batch of null inputs (indicated by the red vertical line in Figure 2) delivers consistent and significant improvement compared to the original LM (although might not be the best improvement). Therefore, for enhancing LM zero-shot performance, we directly adopt the One-batch Calibration as the early stopping criterion.
For few-shot downstream tasks. With the acquisition of a few labeled data, the previous challenge of lacking validation for determining the stopping point in the calibration process is alleviated. We leverage the downstream training dataset constituting samples per class as validation dataset in the calibration.
We take into account above-mentioned empirical findings that for some tasks stopping at one batch of calibration yields optimal LM performance. Relying on the limited size of the validation dataset might fail to identify such stopping points. To this effect, we store both (obtained from one-batch stopping) and (obtained from validation-based stopping) for downstream few-shot leaning tasks. Since is stored in the process of obtaining , this will not result in additional computation overhead. Memory overhead is minimal, as it only requires storing an additional set of updated bias parameters.
@@ 4.1 Generate Null-Meaning Input
We employ null-meaning inputs to probe the intrinsic bias of pre-trained LMs, and then use those bias-reflected outputs to calibrate the LMs. Crafting a diverse set of null-meaning inputs for an averaged output helps prevent overfitting to sub-optimal instances, thereby contributing to the effectiveness of calibration. However, creating numerous null-meaning inputs manually could be laborious and challenging. To enable cost-effective acquisition of various null-meaning data, we utilize GPT-4 API for automatic generation with instructions such as "Please generate null meaning symbols, words, phrases, and sentences, in total <Number>.". Note that this generation process is task-agnostic, ensuring that each generated data contains null information with respect to any downstream task.
@@ 4.2 Select and Build Null-Input Prompt
In-context lrn no demo† | In-context lrn with demo | Prompt FT no demo | Prompt FT with demo | |||||||||
NoCal | OutCal | IntrCal | NoCal | OutCal | IntrCal | NoCal | OutCal | IntrCal | NoCal | OutCal | IntrCal | |
AGNews | ||||||||||||
DBPedia | ||||||||||||
TREC | ||||||||||||
Subj | ||||||||||||
SST-5 | ||||||||||||
Laptop | ||||||||||||
Restaurant | ||||||||||||
@@
To pursue better cohesive integration of the "null" information into the prompts, we prioritize the null-meaning inputs, with which the answer format exhibits higher Next Sentence Prediction (NSP) probability Devlin et al. (2019). Specifically, after we generate a large set of null-meaning inputs and the answer format ans is selected, we employ BERT-large model Devlin et al. (2019) to predict NSP and sort null-meaning inputs by their probabilities. Table 1 shows some generated , with which a specific answer format presents high/low NSP scores. We find this strategy brings advantages in two aspects: (1) It filters some null-meaning inputs that do not conform to normal symbols or text (e.g., "////////////////////"); (2) In the prompts, null-meaning inputs with higher exhibit higher attention scores between the null input and <mask> as shown in Figure 3. This indicates more effective conveyance of the "null" information to the placeholder <mask>, which could facilitate LM deciphering the "null" patterns of the prompts and benefit calibration.
After the sorting, we discard the bottom instances and obtain 800 null-meaning inputs. Specially, for zero-shot downstream tasks, since only one batch of null-meaning inputs is required for calibration in our early-stopping criterion (§ 3.3), we select the from , where is batch size. We show calibration with selection strategy further improves LM performance in § 5.2 Table 3.

### v2

@@ Abstract
Prompt-based learning is susceptible to intrinsic bias present in pre-trained language models (LMs), leading to sub-optimal performance in prompt-based zero/few-shot settings. In this work, we propose a null-input prompting method to calibrate intrinsic bias encoded in pre-trained LMs. Different from prior efforts that address intrinsic bias primarily for social fairness and often involve excessive computational cost, our objective is to explore enhancing LMs’ performance in downstream zero/few-shot learning while emphasizing the efficiency of intrinsic bias calibration. Specifically, we leverage a diverse set of auto-selected null-meaning inputs generated from GPT-4 to probe intrinsic bias of pre-trained LMs. Utilizing the bias-reflected probability distribution, we formulate a distribution disparity loss for bias calibration, where we exclusively update bias parameters ( of total parameters) of LMs towards equal probability distribution. Experimental results show that the calibration promotes an equitable starting point for LMs while preserving language modeling abilities. Across a wide range of datasets, including sentiment analysis and topic classification, our method significantly improves zero/few-shot learning performance of LMs for both in-context learning and prompt-based fine-tuning (on average and , respectively).111Our code is available at https://github.com/kang-ml/prompt_based_bias_calibration.
@@ 1 Introduction
The advent of GPT models Radford et al. (2019); Brown et al. (2020) has catalyzed the transformative prompt-based learning paradigm. The innovative approach of "pre-train, prompt, and predict" Schick and Schütze (2021a); Liu et al. (2023) facilitates fast adaptation of pre-trained language models (LMs) in learning various tasks and empowers LMs’ strong zero/few-shot learning abilities Schick and Schütze (2021b); Gao et al. (2021).
@@
Due to the susceptibility to bias ingrained in pre-trained LMs, prompt-based learning tends to make biased predictions toward some specific answers, thereby impacting performance in prompt-based zero/few-shot settings Zhao et al. (2021); Han et al. (2023). To mitigate this issue and improve LM performance, Zhao et al. (2021) and Holtzman et al. (2022) propose to reweigh LM output probabilities. Han et al. (2023) explores calibrating decision boundaries. While these research has demonstrated substantial improvements, they are primarily designed for in-context learning with frozen pre-trained LMs, leading to two main limitations: (1) They may be not effective in task-specific fine-tuning scenario Jian et al. (2022). Note, however, prompt-based fine-tuning has shown performance improvements over in-context learning Gao et al. (2021); Logan IV et al. (2022). It is particularly important for relatively small-sized LMs. (2) The intrinsic bias encoded in pre-trained LMs persists since these research focuses on output calibration and does not modify LMs.
@@
To address these limitations, we investigate the potential for enhancing the performance of LMs as zero/few-shot learners in classification tasks by calibrating intrinsic bias of pre-trained LMs. This exploration extends to various prompt-based learning scenarios: in-context learning and prompt-based fine-tuning. Prior approaches to mitigate intrinsic bias primarily focus on achieving social fairness, and often require laborious corpora augmentation and costly re-training Huang et al. (2020); Kaneko and Bollegala (2021); Solaiman and Dennison (2021); Li et al. (2023a). To improve efficiency in both data generation and model updates, we propose leveraging auto-generated null-meaning inputs to prompt pre-trained LMs for intrinsic bias probing, and subsequently updating only bias parameters BLM of LMs for bias calibration. Null-meaning inputs are essentially normal text devoid of meaningful content or sentiment. Unlike numerical-zero inputs, they maintain the contextual framework of prompts, ensuring the proper functioning of contextual LMs. Our motivation stems from the expectation that bias-calibrated models should produce uniform probabilities across all categories if the input in a prompt delivers null information Zhao et al. (2021). BLM functions as offsets in neural networks, and strategically updating only BLM could potentially counteract intrinsic bias of pre-trained models, achieving higher efficiency (updating parameters of entire LM). The approach promotes an equitable starting point, and we expect that the light model updates preserve pre-trained models’ language modeling abilities while maintaining the focus on bias calibration, ultimately making LMs better zero/few-shot learners.
@@
The pipeline of our calibration method is illustrated in Figure 1. We use Masked LMs (RoBERTa Liu et al., 2019) for zero/few-shot learning since they generally produce competitive performance in classification tasks and their moderate size facilitates combining prompting with fine-tuning Gao et al. (2021); Liu et al. (2023). First, we utilize GPT-4 API to automatically generate diverse null-meaning inputs including symbols, words, phrases, and sentences. This generation process is downstream task-agnostic. By concatenating each null-meaning input with an answer format ans aligned with the downstream task, we construct null-input prompts (similar to Zhao et al., 2021), e.g., "An empty sentence. It is about <mask>.". For better cohesive integration of the "null" information into the prompts, we additionally devise a filtering strategy to select , to which the answer format ans exhibits relatively strong Next Sentence Prediction (NSP) correlation Devlin et al. (2019). Next, we update BLM with null-input prompts to calibrate intrinsic bias. Given the absence of task-relevant information in these prompts, the anticipated outcome in the parameter updating process is a convergence towards equal output probabilities for each label word. We formulate a customized Kullback–Leibler (KL) divergence loss for gradient descent on BLM to minimize the distribution disparity. Finally, bias-calibrated LMs are applied in downstream prompt-based zero/few-shot learning following Gao et al. (2021).
@@
Extensive experiments on eight classification datasets with four prompt-based learning approaches show that our method significantly improves LMs’ zero/few-shot performance, and outperforms output-calibration methods.
2 Related Work
Impact of intrinsic bias on downstream LM performance. Intrinsic bias in pre-trained LMs stems from imbalances present in extensive pre-training corpora. Higher frequency of specific terms in those corpora could lead to common token bias Zhao et al. (2021). Additionally, frequent co-occurrence of certain terms with specific sentiment in pre-training could introduce association bias Cao et al. (2022). Because of those intrinsic bias, prompt-based predictions by pre-trained LMs are prone to bias towards some specific answers, resulting in sub-optimal performance in downstream tasks Zhao et al. (2021); Han et al. (2023).
Mitigating strategies. Research has focused on counteracting the bias solely at the output prediction stage, without modifying pre-trained LMs. For example, Zhao et al. (2021) introduces contextual calibration and Holtzman et al. (2022) presents Domain Conditional Pointwise Mutual Information to reweigh answer scores. Min et al. (2022) explores computing the probability of the input conditioned on the label. Han et al. (2023) proposes to calibrate decision boundaries. However, these studies mainly demonstrate their effectiveness for in-context learning using frozen pre-trained LMs, without addressing the intrinsic bias encoded in the LMs. Other research on mitigating intrinsic bias primarily targets removing social bias Dinan et al. (2020); Huang et al. (2020); Cheng et al. (2021); Zhou et al. (2023), often employing costly data augmentation and re-training, and as a by-product, degrades language modeling abilities Meade et al. (2022).
Efficiently calibrating intrinsic bias in pre-trained LMs for enhancing downstream zero/few-shot learning performance is an open research problem. We introduce a parameter-efficient intrinsic-bias calibration method leveraging automatically constructed null-input prompts, which significantly improves zero/few-shot learning of LMs.
Parameter-efficient fine-tuning (PEFT) for downstream tasks. It has been demonstrated that fine-tuning a very small portion of model parameters can achieve performance on par with fine-tuning the entire set of parameters. People propose integrating small, trainable adapter modules between model layers Bapna and Firat (2019); Houlsby et al. (2019), coupled with further optimization using low-rank adaptations (LoRA) Hu et al. (2021). Some other research focuses on prompt tuning Lester et al. (2021); Li and Liang (2021); Gu et al. (2022); Guo et al. (2022) which only tunes continuous prompt embeddings for efficiently adapting pre-trained LMs to downstream tasks.
Our method provides a unique perspective of enhancing LM performance on downstream tasks through efficient intrinsic-bias calibration. We update only bias parameters of pre-trained LMs with null-input prompts in calibration. Contrary to adapters and LoRA which would need sufficient labeled data to learn new matrices, we do not introduce new matrices to pre-trained LMs, preserving LMs’ few-shot learning capabilities. Moreover, our approach does not necessarily require target-domain data (whether labeled or unlabeled), enabling fully unsupervised deployment, particularly advantageous for zero-shot setting.
@@ 3.1 Task Formulation
where denotes uniform probability distribution and represents the simplified form of .
@@ 3.2 Update Only Bias Parameters
While intrinsic bias may be encoded across various parts of pre-trained LMs, one question arises: is it essential to update the entire model, or is there a more efficient alternative that can achieve comparable effectiveness in intrinsic bias calibration? We propose to only update bias parameters BLM, with the following rationale: (i) BLM constitutes less than of total LM parameters, offering significant memory and computation cost saving compared to updating entire LM. (ii) Weight parameters WLM222WLM also includes embedding parameters in our context. may carry crucial pre-existing knowledge for language modeling, which risks impairment with a full model update Meade et al. (2022). BLM, often overlooked in LM research, serves as offsets in neural network layers. Strategic updates may counteract intrinsic bias while potentially preserving language modeling abilities. (iii) Empirical research on efficient fine-tuning has demonstrated the important role of bias parameters in LMs Ben Zaken et al. (2022); Logan IV et al. (2022).
@@ 3.3 Early Stopping of Calibration
We aim to obtain LM with improved zero/few-shot performance at the calibration stopping point. An overly calibrated model may simply produce uniform probability predictions regardless of input information. To avoid this, we develop specialized early stopping strategies depending on whether the downstream task is zero-shot or few-shot.
For zero-shot downstream tasks. Determining the calibration stopping point for optimal zero-shot learning performance is challenging due to the absence of labeled data for validation during calibration. To discern the patterns of a good stopping point, we first conduct empirical experiments by validating LM zero-shot performance on the entire test dataset after each calibration batch (consisting of null-meaning inputs) across different calibration learning rates (Figure 7 in Appendix A). As shown in Figure 2, with optimal calibration learning rate, model performance exhibits significant improvements in the first one/few calibration batches with low variance, and then starts to degrade and becomes unstable. The low performance and instability at the calibration tail confirm our assumption on the detrimental effects of excessive calibration on LM’s modeling abilities. Notably, calibration with only one batch of null inputs (indicated by the red vertical line in Figure 2) delivers consistent and significant improvement compared to the original LM (although might not be the best improvement). Therefore, for enhancing LM zero-shot performance, we directly adopt the One-batch Calibration as the early stopping criterion.
For few-shot downstream tasks. With the acquisition of a few labeled downstream data, the previous challenge of lacking validation for determining the stopping point in the calibration process is alleviated. We utilize the small amount of labeled data as validation dataset to set a stopping criterion for calibration. Additionally, we take into account above-mentioned empirical findings that, for some tasks, stopping at one batch of calibration yields optimal LM performance. Relying on the limited size of might fail to identify such stopping points. To this effect, we store both (obtained from one-batch stopping) and (obtained from validation-based stopping) for downstream few-shot leaning tasks. Since is stored in the process of obtaining , this will not result in additional computation overhead. Memory overhead is minimal, as it only requires storing an additional set of updated bias parameters.
@@ 4.1 Generate Null-Meaning Input
We employ null-meaning inputs to probe the intrinsic bias of pre-trained LMs, and then use those bias-reflected outputs to calibrate the LMs. Crafting a diverse set of null-meaning inputs for an averaged output helps prevent overfitting to sub-optimal instances, thereby contributing to the effectiveness of calibration. To enable cost-effective acquisition of various null-meaning data, we utilize GPT-4 API for automatic generation with instructions such as "Please generate null meaning symbols, words, phrases, and sentences, in total <Number>.". This process is task-agnostic, generating data that contains null information with respect to any downstream task. Note that null information is not equivalent to neutral sentiment, as it carries no inherent meaning or contextual sentiment implications. We further validate this through t-SNE van der Maaten and Hinton (2008) visualization in Appendix A Figure 6.
@@ 4.2 Select and Build Null-Input Prompt
In-context lrn no demo† | In-context lrn with demo | Prompt FT no demo | Prompt FT with demo | ||||||||||||
NoCal | OutCal | IntrCal | NoCal | OutCal | IntrCal | NoCal | OutCal | IntrCal | NoCal | OutCal | IntrCal | ||||
AGNews | |||||||||||||||
DBPedia | |||||||||||||||
TREC | |||||||||||||||
Subj | |||||||||||||||
SST-5 | |||||||||||||||
Laptop | |||||||||||||||
Restaurant | |||||||||||||||
@@
To pursue better cohesive integration of the "null" information into the prompts, we prioritize the null-meaning inputs, with which the answer format exhibits higher Next Sentence Prediction (NSP) probability Devlin et al. (2019). Specifically, after we generate a large set of null-meaning inputs and the answer format ans is selected, we employ BERT-large model Devlin et al. (2019) to predict NSP and sort null-meaning inputs by their probabilities. Table 1 shows some generated , with which a specific answer format presents high/low NSP scores. After the sorting, we retain the top instances (800 in total), which maintains the diversity among the selected samples. We observe that null inputs with low NSP scores are typically randomly-combined alphabet letters and symbols. These samples may have minimal occurrences in pre-training corpora. The low NSP scores can be attributed to RoBERTa’s lack of comprehension of their meanings in context. Their representations extracted by LM might have high variance, which might impact the stability and effectiveness of calibration. We show calibration with the selection strategy further improves LM performance in § 5.2 Table 3.